The following statements are for R Company: R Company Income Statement Net Sales Cost of Goods Sold Gross Profit Operating Expenses Net Income $400,000 Percent 240,000 160,000 100,000 60,000 R Company Balance Sheet 30,000 CCCC Cash Land Total Assets 270,000 $ 300,000 $120,000 DDDD $180,000 Total Liabilities Total Equity
Applying vertical analysis find the value for AAAA, BBBB, CCCC, DDDD.
Answer #1

Answer

  • Based on Vertical Analysis:

Calculation of AAAA

= (Operating Expenses / Net Sales ) x 100

= (100000/400000) x 100 = 25.00%

Calculation of BBBB

= (Net Income/Net sales) x 100

= ( 60000/400000) x 100

= 15.00%

Calculation of CCCC

= (Cash/Total Assets) x 100

= (30000/300000) x 100

= 10.00%

Calculation of DDDD

= [Total Liabilities/ (Total Liabilities + Total Equity)] x 100

= [120000 / (120000 + 180000)] x 100

= [120000 / 300000] x 100

= 40.00%